At 25°C, what is the hydroxide ion concentration, [OH−], in an aqueous solution with a hydrogen ion concentration of [H+]=2.7×10−7 M?
1
Expert's answer
2019-07-31T08:15:54-0400

According to the pH concept: pH + pOH = 14, pH = -log[H+], and pOH = -log[OH-].

As a result:

-log[H+] + -log[OH-] = 14

[OH-] = (1×10-14) / [H+]

[OH-] = (1×10-14) / 2.7×10−7 M = 0.37×10−7 M

Answer: 0.37×10−7 M